### v2

# Computing Evolutionarily Stable Strategies in Imperfect-Information Games

Abstract
We present an algorithm for computing evolutionarily stable strategies (ESSs) in symmetric perfect-recall extensive-form games of imperfect information. Our main algorithm is for two-player games, and we describe how it can be extended to multiplayer games. The algorithm is sound and computes all ESSs in nondegenerate games and a subset of them in degenerate games which contain an infinite continuum of symmetric Nash equilibria. The algorithm is anytime and can be stopped early to find one or more ESSs. We experiment on an imperfect-information cancer signaling game as well as random games to demonstrate scalability.
1 Introduction
The standard solution concept for evolutionary game theory, which typically models strategic interactions between evolving biological populations, is evolutionarily stable strategy (ESS). A mixed strategy in a two-player symmetric game is an evolutionarily stable strategy if it is robust to being overtaken by a mutation strategy. Formally, mixed strategy $\mathbf{x}^{\star}$ is an ESS if for every mixed strategy $\mathbf{x}$ that differs from $\mathbf{x}^{\star}$, there exists $\varepsilon_{0}=\varepsilon_{0}(\mathbf{x})>0$ such that, for all $\varepsilon\in(0,\varepsilon_{0})$,
$(1-\varepsilon)u_{1}(\mathbf{x},\mathbf{x}^{\star})+\varepsilon u_{1}(\mathbf{x},\mathbf{x})<(1-\varepsilon)u_{1}(\mathbf{x}^{\star},\mathbf{x}^{\star})+\varepsilon u_{1}(\mathbf{x}^{\star},\mathbf{x}).$
From a biological perspective, we can interpret $\mathbf{x}^{\star}$ as a distribution among “normal” individuals within a population, and consider a mutation that makes use of strategy $\mathbf{x}$, assuming that the proportion of the mutation in the population is $\varepsilon$. In an ESS, the expected payoff of the mutation is smaller than the expected payoff of a normal individual, and hence the proportion of mutations will decrease and eventually disappear over time, with the composition of the population returning to being mostly $\mathbf{x}^{\star}$. An ESS is therefore a mixed strategy of the column player that is immune to being overtaken by mutations. ESS was initially proposed by mathematical biologists motivated by applications such as population dynamics (e.g., maintaining robustness to mutations within a population of humans or animals) Maynard Smith and Price (1973); Maynard Smith (1982). A common example game is the 2x2 game where strategies correspond to an “aggressive” Hawk or a “peaceful” Dove strategy. A paper has recently proposed a similar game in which an aggressive malignant cell competes with a passive normal cell for biological energy, which has applications to cancer eradication Dingli et al. (2009).
A substantial body of work has explored evolutionary game dynamics in extensive-form strategic settings, beginning with the foundational analyses of Selten on evolutionarily stable strategies in multi-stage games with imperfect information. Selten formalized how the ESS concept extends beyond normal form to behavioral strategies in extensive-form games, and showed that mutation stability can depend on sequential information sets and continuation payoffs rather than only aggregate action frequencies Selten (1983, 1988). Subsequent work by Cressman provided a systematic treatment of evolutionary dynamics in extensive-form games, developing dynamical systems and stability results for behavioral strategies in games with sequential structure, chance nodes, and imperfect information Cressman (2003). This line of research established that sequential aspects of play—specifically the information available at each decision point—can fundamentally alter evolutionary stability conditions compared to normal-form representations.
Additional contributions have examined evolutionary dynamics on extensive-form or multi-stage games from both theoretical and computational perspectives. Matros analyzed replicator-type dynamics defined directly on the sequence form of an extensive-form game Matros (2001), while Gatti et al. studied efficient computation of evolutionary trajectories in EFGs and demonstrated that evolutionary behavior can differ sharply from normal-form reductions even in small games Gatti et al. (2013). Beyond these general frameworks, many evolutionary models naturally incorporate imperfect information through signaling or hidden-type structures. Evolutionarily stable behavior in signaling games has been extensively investigated—for example, Jäger’s ESS conditions for costly signaling equilibria Jäger (2008) and Huttegger’s evolutionary analysis of Lewis signaling systems Huttegger (2010)—all of which rely on imperfect information at the level of types or messages. Related work on Bayesian evolutionary games similarly treats agents’ types or preferences as private information, yielding evolutionarily stable outcomes in settings with incomplete information such as cooperation under hidden types Akçay and Cleve (2012) and preference evolution in Bayesian environments Alger and Weibull (2013).
Together, this literature demonstrates both the importance and the difficulty of analyzing evolutionary stability in extensive-form settings: sequential play, hidden information, and structured strategy spaces significantly complicate the evaluation of mutant strategies. However, despite substantial theoretical work, there is currently no general algorithmic framework for computing ESS in extensive-form games, even though such games arise naturally in evolutionary models of signaling, conflict escalation, communication, sequential cooperation, and biological or ecological interactions. This motivates the present work, which develops the first practical algorithm for identifying evolutionarily stable strategies in general extensive-form games of imperfect information.
While Nash equilibrium is defined for general multiplayer games, ESS is traditionally defined specifically for two-player symmetric games. ESS is a refinement of Nash equilibrium. In particular, if $\mathbf{x}^{\star}$ is an ESS, then $(\mathbf{x}^{\star},\mathbf{x}^{\star})$ (i.e., the strategy profile where both players play $\mathbf{x}^{\star}$) is a (symmetric) Nash equilibrium Maschler et al. (2013). Of course the converse is not necessarily true (not every symmetric Nash equilibrium is an ESS), or else ESS would be a trivial refinement. In fact, ESS is not guaranteed to exist in games with more than two pure strategies per player (while Nash equilibrium is guaranteed to exist in all finite games). For example, while rock-paper-scissors has a mixed strategy Nash equilibrium (which puts equal weight on all three actions), it has no ESS Maschler et al. (2013) (that work considers a version where payoffs are 1 for a victory, 0 for loss, and $\frac{2}{3}$ for a tie).
There exists a polynomial-time algorithm for computing Nash equilibrium (NE) in two-player zero-sum normal-form games, while for two-player non-zero-sum and multiplayer games computing an NE is PPAD-complete and it is widely conjectured that no efficient (polynomial-time) algorithm exists. However, several algorithms have been devised that perform well in practice. The problem of computing whether a game has an ESS was shown to be both NP-hard and CO-NP hard and also to be contained in $\Sigma^{P}_{2}$ (the class of decision problems that can be solved in nondeterministic polynomial time given access to an NP oracle) Etessami and Lochbihler (2008). Subsequently it was shown that the exact complexity of this problem is that it is $\Sigma^{P}_{2}$-complete Conitzer (2013), and even more challenging for more than two players Blanc and Hansen (2021). Note that this result is for determining whether an ESS exists (as discussed above there exist games which have no ESS), not for the complexity of computing an ESS in games for which one exists. Thus, computing an ESS is significantly more difficult than computing an NE, which is not surprising since it is a refinement. Several approaches have been proposed for computing ESS in two-player normal-form games Haigh (1975); Abakuks (1980); Broom and Rychtář (2013); Bomze (1992); McNamara et al. (1997), and recently an algorithm has been developed for multiplayer normal-form games Ganzfried (2025a). However, these algorithms are based on support enumeration and are not applicable to extensive-form games.
2 Evolutionarily Stable Strategies
We first review the standard definition of evolutionarily stable strategies for symmetric normal-form games then describe how the definition can be extended to extensive-form games. A normal-form game consists of a finite set of players $N=\{1,\ldots,n\}$, a finite set of pure strategies $S_{i}$ for each player $i$, and a real-valued utility for each player for each strategy vector (aka strategy profile), $u_{i}:\times_{i}S_{i}\rightarrow\mathbb{R}$. In a symmetric normal-form game, all strategy spaces $S_{i}$ are equal and the utility functions satisfy the following symmetry condition: for every player $i\in N$, pure strategy profile $(s_{1},\ldots,s_{n})\in S^{n}$, and permutation $\pi$ of the players,
$u_{i}(s_{1},\ldots,s_{n})=u_{\pi(i)}(s_{\pi(1)},\ldots,s_{\pi(n)}).$
This allows us to remove the player index of the utility function and just write $u(s_{1},\ldots,s_{n}),$ where it is implied that the utility is for player 1 (we can simply permute the players to obtain the utilities of the other players). We will still write $u_{i}$ for notational convenience, but note that only a single utility function must be specified which will apply to all players.
Let $\Sigma_{i}$ denote the set of mixed strategies of player $i$ (probability distributions over elements of $S_{i}$). If players follow mixed strategy profile $\mathbf{x}=(\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(n)}),$ where $\mathbf{x}^{(i)}\in\Sigma_{i}$, the expected payoff to player $i$ is
$u_{i}(\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(n)})=\sum_{s_{1},\ldots,s_{n}\in S}x^{(1)}_{s_{1}}\cdots x^{(n)}_{s_{n}}u_{i}(s_{1},\ldots,s_{n}).$
We write $u_{i}(\mathbf{x})=u_{i}(\mathbf{x}^{(i)},\mathbf{x}^{(-i)})$, where $\mathbf{x}^{(-i)}$ denotes the vector of strategies of all players except $i$. If all players follow the same mixed strategy $\mathbf{x}$, then for all players $i$ we have
$u_{i}(\mathbf{x})=u_{i}(\mathbf{x},\ldots,\mathbf{x})=\sum_{s_{1},\ldots,s_{n}\in S}x_{s_{1}}\cdots x_{s_{n}}u_{i}(s_{1},\ldots,s_{n}).$
Definition 1.
A mixed strategy profile $\mathbf{x}^{\star}$ is a Nash equilibrium if for each player $i\in N$ and for each mixed strategy $\mathbf{x}^{(i)}\in\Sigma_{i}$: $u_{i}(\mathbf{x}^{*(i)},\mathbf{x}^{*(-i)})\geq u_{i}(\mathbf{x}^{(i)},\mathbf{x}^{*(-i)}).$
Definition 2.
A mixed strategy profile $\mathbf{x}^{\star}$ in a symmetric normal-form game is a symmetric Nash equilibrium if it is a Nash equilibrium and: $\mathbf{x}^{*(1)}=\mathbf{x}^{*(2)}=\ldots=\mathbf{x}^{*(n)}.$
Definition 3.
A mixed strategy $\mathbf{x}^{\star}\in\Sigma_{1}$ is evolutionarily stable in a symmetric normal-form game if for each mixed strategy $\mathbf{x}\neq\mathbf{x}^{\star}$ exactly one of the following conditions holds:
  1. 1.
    $u_{1}(\mathbf{x}^{*},\mathbf{x}^{*},\ldots,\mathbf{x}^{*})>u_{1}(\mathbf{x},\mathbf{x}^{*},\ldots,\mathbf{x}^{*})$,
  2. 2.
    $u_{1}(\mathbf{x}^{*},\mathbf{x}^{*},\ldots,\mathbf{x}^{*})=u_{1}(\mathbf{x},\mathbf{x}^{*},\ldots,\mathbf{x}^{*})$ and $u_{1}(\mathbf{x}^{*},\mathbf{x},\mathbf{x}^{*},\ldots,\mathbf{x}^{*})>u_{1}(\mathbf{x},\mathbf{x},\mathbf{x}^{*},\ldots,\mathbf{x}^{*})$.
It has been proven that every symmetric normal-form game has at least one symmetric Nash equilibrium Nash (1951). It is clear from Definition 3 that every evolutionarily stable strategy in a symmetric normal-form game must be a symmetric Nash equilibrium (SNE). Thus, a natural approach for ESS computation is to first compute SNE and then perform subsequent procedures to determine whether they are ESS.
Imperfect-information games are modeled using extensive-form game trees, where play proceeds from the root node to a terminal leaf node at which point all players receive payoffs. Each non-terminal node has an associated player (possibly chance) that makes the decision at that node. These nodes are partitioned into information sets, where the player whose turn it is to move cannot distinguish among the states in the same information set. Therefore, in any given information set, a player must choose actions with the same distribution at each state contained in the information set. If no player forgets information that they previously knew, we say that the game has perfect recall. A (mixed) strategy for player $i,$ $\sigma_{i}\in\Sigma_{i},$ is a function that assigns a probability distribution over all actions at each information set belonging to $i$.
In order to present our new approach we review the sequence-form representation for two-player extensive-form games of imperfect information Koller et al. (1994). Rather than operate on the full pure strategy space, which has size exponential in the size of the game tree, the sequence-form works with sequences of actions along trajectories from the root node to leaf nodes. For player 1, the matrix $\mathbf{E}$ is defined where each row corresponds to an information set (including an initial row for the “empty” information set), and each column corresponds to an action sequence (including an initial row for the “empty” action sequence). In the first row of $\mathbf{E}$ the first element is 1 and all other elements are 0; subsequent rows have -1 for the entries corresponding to the action sequence leading to the root of the information set, and 1 for all actions that can be taken at the information set (and 0 otherwise). Thus $\mathbf{E}$ has dimension $c_{1}\times d_{1}$, where $c_{i}$ is the number of information sets for player $i$ and $d_{i}$ is the number of action sequences for player $i$. Matrix $\mathbf{F}$ is defined analogously for player 2. The vector $\mathbf{e}$ is defined to be a column vector of length $c_{1}$ with 1 in the first position and 0 in other entries, and vector $\mathbf{f}$ is defined with length $c_{2}$ analogously. The matrix $\mathbf{A}$ is defined with dimension $d_{1}\times d_{2}$ where entry $A_{ij}$ gives the payoff for player 1 when player 1 plays action sequence $i$ and player 2 plays action sequence $j$ multiplied by the probabilities of chance moves along the path of play. The matrix $\mathbf{B}$ of player 2’s payoffs is defined analogously.
For two-player non-zero-sum games, the problem of finding a Nash equilibrium is the feasibility problem of finding $\mathbf{x},$ $\mathbf{y},$ $\mathbf{p},$ $\mathbf{q}$, such that Koller et al. (1994):
$-\mathbf{A}\mathbf{y}+\mathbf{E}^{\top}\mathbf{p}$ $\geq\mathbf{0}$ (1)
$-\mathbf{B}^{\top}\mathbf{x}+\mathbf{F}^{\top}\mathbf{q}$ $\geq\mathbf{0}$
$\mathbf{E}\mathbf{x}$ $=\mathbf{e}$
$\mathbf{F}\mathbf{y}$ $=\mathbf{f}$
$\mathbf{x}$ $\geq\mathbf{0}$
$\mathbf{y}$ $\geq\mathbf{0}$
$\mathbf{x}^{\top}(-\mathbf{A}\mathbf{y}+\mathbf{E}^{\top}\mathbf{p})$ $=0$
$\mathbf{y}^{\top}(-\mathbf{B}\mathbf{x}+\mathbf{F}^{\top}\mathbf{q})$ $=0$
The final two constraints are called complementarity slackness conditions, and the full system is known as a linear complementarity problem (LCP). This LCP can be solved using Lemke’s algorithm Lemke (1965) or the related Lemke-Howson algorithm Lemke and Howson (1964).
A two-player extensive-form game is symmetric if the two players’ game trees are isomorphic, including identical information sets and available actions at corresponding nodes; if all chance moves and their probabilities coincide under this isomorphism; and if the payoff functions satisfy
$u_{1}(z)=u_{2}(\pi(z))$
for every terminal history $z$, where $\pi$ is the player-swapping automorphism of the game tree. Under symmetry, we may identify a behavioral strategy with its realization plan $\mathbf{x}$ in sequence form, subject to the information-set consistency constraints $\mathbf{E}\mathbf{x}=\mathbf{e}$, $\mathbf{x}\geq\mathbf{0}.$ Let $\boldsymbol{\sigma}$ denote a behavioral strategy, and let $u(\boldsymbol{\sigma},\boldsymbol{\tau})$ be the expected payoff of playing $\boldsymbol{\sigma}$ against an opponent playing $\boldsymbol{\tau}$. Because the game is symmetric, the classical definition of evolutionary stability applies directly to behavioral strategies, with $\mathbf{B}=\mathbf{A}^{\top},$ $\mathbf{F}=\mathbf{E},$ $\mathbf{f}=\mathbf{e}.$
Definition 4.
A behavioral strategy $\boldsymbol{\sigma}^{\star}$ is an evolutionarily stable strategy in a symmetric perfect-recall extensive-form imperfect-information game if for every behavioral strategy $\boldsymbol{\sigma}\neq\boldsymbol{\sigma}^{\star}$, exactly one of the following holds:
  1. 1.
    $u(\boldsymbol{\sigma}^{\star},\boldsymbol{\sigma}^{\star})>u(\boldsymbol{\sigma},\boldsymbol{\sigma}^{\star})$, or
  2. 2.
    $u(\boldsymbol{\sigma}^{\star},\boldsymbol{\sigma}^{\star})=u(\boldsymbol{\sigma},\boldsymbol{\sigma}^{\star})$ and $u(\boldsymbol{\sigma}^{\star},\boldsymbol{\sigma})>u(\boldsymbol{\sigma},\boldsymbol{\sigma})$.
Evolutionary stability therefore requires that $\boldsymbol{\sigma}^{\star}$ perform strictly better than any mutant strategy $\boldsymbol{\sigma}$ when both are played against the incumbent population, and—when the two strategies perform equally well against the incumbent—that $\boldsymbol{\sigma}^{\star}$ outperform $\boldsymbol{\sigma}$ in direct competition. In the sequence-form representation, any candidate behavioral strategy must satisfy the linear constraints $\mathbf{E}\mathbf{x}=\mathbf{e}$ and $\mathbf{x}\geq\mathbf{0}$, ensuring that evolutionary comparisons are carried out over feasible realization plans within the extensive-form structure.
Our algorithm will focus on the setting of two-player symmetric perfect-recall extensive-form games. We note that sequence-form representation extends naturally to $n$-player games, as do the definitions of symmetric games and evolutionarily stable strategies. We will briefly discuss how our algorithm can be extended to apply to $n$ players for $n>2.$
3 Algorithm
Our main algorithm for ESS computation in two-player symmetric perfect-recall extensive-form games is given in Algorithm 1. Our method is an anytime algorithm that iteratively searches for a new SNE $\mathbf{x},$ and then tests whether $\mathbf{x}$ is an ESS. We can halt the algorithm as soon as one ESS is found, or continue it to find as many ESSs as possible until an arbitrary termination criterion is found. The algorithm takes the sequence-form game matrices $(\mathbf{E},\mathbf{e},\mathbf{A})$ as input, and uses several hyperparameters $(\delta,\varepsilon_{p},\varepsilon_{s})$ for numerical precision. The values used in the experiments are given in Table 1.
Algorithm 1 first computes an initial symmetric Nash equilibrium $\mathbf{x}$ by running the procedure given in Algorithm 2. From Nash’s theorem we know that a symmetric Nash equilibrium always exists. We then add $\mathbf{x}$ to the list of known SNE and calculate the expected payoff of playing $\mathbf{x}$ against itself: $v^{\star}=\mathbf{x}^{\top}\mathbf{A}\mathbf{x}.$ We then run Algorithm 4 to test whether $\mathbf{x}$ satisfies the ESS conditions. We then run Algorithm 3 to compute a new undiscovered SNE (if one exists), and repeat this procedure until a desired termination criterion is met.
Algorithm 1 ESS search in symmetric two-player extensive-form game of imperfect information
0: Sequence-form game matrices $(\mathbf{E},\mathbf{e},\mathbf{A})$, parameters $\delta>0$, $\varepsilon_{p}>0$, $\varepsilon_{s}>0$, termination criterion
0: List $\mathrm{ESS\_list}$
[…]
7:if $\text{status}=\text{INFEASIBLE}$ or $F^{\star}>\varepsilon_{p}$ then
8:  $\mathrm{ESS\_list}\leftarrow\mathrm{ESS\_list}\cup\{\mathbf{x}\}$
9:end if
10:while termination criterion not met do
11:  $(\mathbf{x},t^{\star})\leftarrow\textsc{NEW\_SNE}(\mathbf{E},\mathbf{e},\mathbf{A},\mathrm{SNE\_list})$
12:  if $t^{\star}<\varepsilon_{s}$ then
13:   break
14:  end if
15:  $\mathrm{SNE\_list}\leftarrow\mathrm{SNE\_list}\cup\{\mathbf{x}\}$
16:  $v^{\star}\leftarrow\mathbf{x}^{\top}\mathbf{A}\mathbf{x}$
[…]
18:  if $\text{status}=\text{INFEASIBLE}$ or $F^{\star}>\varepsilon_{p}$ then
19:   $\mathrm{ESS\_list}\leftarrow\mathrm{ESS\_list}\cup\{\mathbf{x}\}$
20:  end if
21:end while
22:return $\mathrm{ESS\_list}$
Algorithm 2 computes the initial SNE by solving a quadratically-constrained feasibility program that is a simpler version of the LCP for computing a Nash equilibrium in general (potentially asymmetric) games given by equation system (1). Algorithm 3 attempts to compute a subsequent SNE given a list of the SNE that were previously obtained. Let $\{\mathbf{x}^{(1)},\ldots,\mathbf{x}^{(T)}\}$ denote current list of SNE. The algorithm finds an SNE that satisfies the same constraints as those in Algorithm 2 and maximizes the minimum (squared) L2 distance from an existing SNE. It does this by maximizing a new scalar variable $t$ with additional constraints
$\|\mathbf{x}-\mathbf{x}^{(j)}\|_{2}^{2}\geq t\mbox{ for all }j=1,\ldots,T.$
This quadratically-constrained feasibility program will always be feasible, since any of the previously found SNE $\mathbf{x}^{(i)}$ is a feasible solution (with optimal objective $t^{\star}=0$). Algorithm 1 then tests whether the newly found candidate SNE $\mathbf{x}$ is numerically distinct from the existing SNE by comparing $t^{\star}$ to the numerical separation parameter $\varepsilon_{s}.$ If $t^{\star}<\varepsilon_{s}$ then we conclude that no additional SNE exist that are numerically distinct from the existing SNE set, and halt the algorithm. Note that in degenerate games where a continuum of SNE exist our algorithm may fail to find all SNE (and also all ESSs), though in such a case any algorithm would fail to enumerate all SNE and/or ESSs. Therefore our algorithm is technically incomplete and not guaranteed to find all ESS, though it is sound in that all ESSs found are actually ESSs (i.e., the algorithm is guaranteed to output a subset all ESSs).
[…]
1: Find $\mathbf{x},\mathbf{p},\mathbf{s}$ such that:
2:   $\mathbf{E}\mathbf{x}=\mathbf{e}$,  $\mathbf{x}\geq\mathbf{0}$
3:   $\mathbf{E}^{\top}\mathbf{p}-\mathbf{A}\mathbf{x}-\mathbf{s}=\mathbf{0}$,  $\mathbf{s}\geq\mathbf{0}$
4:   $x_{i}s_{i}=0$ for all $i$
5:return $\mathbf{x}$
Algorithm 3 NEW_SNE $(\mathbf{E},\mathbf{e},\mathbf{A},\mathrm{SNE\_list})$
[…]
1: Maximize $t$ over variables $\mathbf{x},\mathbf{p},\mathbf{s},t$ subject to:
2:   $\mathbf{E}\mathbf{x}=\mathbf{e}$,  $\mathbf{x}\geq\mathbf{0}$
3:   $\mathbf{E}^{\top}\mathbf{p}-\mathbf{A}\mathbf{x}-\mathbf{s}=\mathbf{0}$,  $\mathbf{s}\geq\mathbf{0}$
4:   $x_{i}s_{i}=0$ for all $i$
5:   $\|\mathbf{x}-\mathbf{x}^{(j)}\|_{2}^{2}\geq t$ for all $j=1,\ldots,T$
[…]
7:return $(\mathbf{x}_{\text{new}},t^{\star})$
Given a newly found SNE $\mathbf{x}$ and its payoff against itself $v^{\star}$, Algorithm 4 tests whether $\mathbf{x}$ is in fact an ESS. Since $\mathbf{x}$ is an SNE, we know that $\mathbf{x}^{\top}\mathbf{A}\mathbf{x}\geq\mathbf{y}^{\top}\mathbf{A}\mathbf{x}$ for all realization sequences $\mathbf{y}.$ We now need to test whether any mutants $\mathbf{y}$ exist that violate both conditions of Definition 4. Note that we cannot have $\mathbf{y}^{\top}\mathbf{A}\mathbf{x}>\mathbf{x}^{\top}\mathbf{A}\mathbf{x}$ since that would violate the condition that $\mathbf{x}$ is an SNE. If $\mathbf{x}^{\top}\mathbf{A}\mathbf{x}>\mathbf{y}^{\top}\mathbf{A}\mathbf{x}$, then the mutant fails to violate the first condition. So the only way that a mutant can successfully violate both conditions is if both $\mathbf{x}^{\top}\mathbf{A}\mathbf{x}=\mathbf{y}^{\top}\mathbf{A}\mathbf{x}$ and $\mathbf{x}^{\top}\mathbf{A}\mathbf{y}\leq\mathbf{y}^{\top}\mathbf{A}\mathbf{y}.$ To test this, we will minimize the objective $F(\mathbf{y})=\mathbf{x}^{\top}\mathbf{A}\mathbf{y}-\mathbf{y}^{\top}\mathbf{A}\mathbf{y},$ subject to the constraint $\mathbf{y}^{\top}\mathbf{A}\mathbf{x}=v^{\star},$ where we have already computed $v^{\star}=\mathbf{x}^{\top}\mathbf{A}\mathbf{x}.$ To ensure that $\mathbf{y}$ is numerically distinct we impose the additional constraint $\|\mathbf{y}-\mathbf{x}\|_{2}^{2}\geq\delta^{2}$. If the optimal objective value $F^{\star}$ is less than or equal to zero then there exists a mutant that violates the conditions and $\mathbf{x}$ is not an ESS: otherwise $\mathbf{x}$ is an ESS. To test this in practice numerically Algorithm 1 compares $F^{\star}$ to $\varepsilon_{p}.$ It is also worth noting that if the problem is infeasible, this means that there do not exist any mutants numerically distinct from $\mathbf{x}$ that violate the ESS conditions; in this case we can also conclude that $\mathbf{x}$ is an ESS.
[…]
1: Minimize $F(\mathbf{y})=\mathbf{x}^{\top}\mathbf{A}\mathbf{y}-\mathbf{y}^{\top}\mathbf{A}\mathbf{y}$ subject to:
2:   $\mathbf{E}\mathbf{y}=\mathbf{e}$,  $\mathbf{y}\geq\mathbf{0}$
3:   $\mathbf{y}^{\top}\mathbf{A}\mathbf{x}=v^{\star}$
4:   $\|\mathbf{y}-\mathbf{x}\|_{2}^{2}\geq\delta^{2}$
5:if infeasible then
6:  return (INFEASIBLE, $\bot$)
7:end if
[…]
9:return (FEASIBLE, $F^{\star}$)
Parameter Value
$\varepsilon_{s}$ $10^{-3}$
$\varepsilon_{p}$ $10^{-5}$
$\delta$ $10^{-2}$
Table 1: Parameter values used in the algorithm.
4 Experiments
To solve the subproblems given in Algorithms 2, 3, and 4 numerically, we used the nonconvex quadratic solver from Gurobi version 13.0.0, which guarantees global optimality Gurobi Optimization, LLC (2025), with Java version 14.0.2. The problem in Algorithms 2 is a quadratically-constrained feasibility program (QCFP), the problem in Algorithm 3 is a quadratically constrained program (QCP), and the problem in Algorithm 4 is a quadratically-constrained quadratic program (QCQP), all of which are nonconvex. We used an Intel Core i7-1065G7 processor with base clock speed of 1.30 GHz (and maximum turbo boost speed of up to 3.9 GHz) with 16 GB of RAM under 64-bit Windows 11 (8 logical cores/threads). We used the numerical parameter values given in Table 1. We used the default feasibility tolerance in Gurobi of $10^{-6}$ (meaning that it is possible that constraints are violated by up to $10^{-6}$). Thus our numerical parameters $\varepsilon_{s}$, $\varepsilon_{p}$, and $\delta^{2}$ are all well above this value. We also performed a numerical procedure to the SNE candidates $\mathbf{x}$ returned by Algorithms 2 and 3 that set all action sequence values below $\varepsilon_{CLIP}=10^{-5}$ to zero and renormalized. For the termination criterion we halt the algorithm if $t^{\star}<\varepsilon_{s}.$
We first evaluate our algorithm on a biologically motivated two-player symmetric imperfect-information game that models clonal interaction under uncertain therapy conditions. At the root, a chance node selects the systemic therapy intensity $T\in\{\text{Low},\text{High}\}$, which is not observed by the players. Each intensity is selected with probability 0.5. Each clone $i\in\{1,2\}$ then receives a private noisy signal $S_{i}\in\{\text{fav},\text{unfav}\}$, where favorable signals are more likely when therapy is low and unfavorable signals are more likely when therapy is high. Specifically, we assume that each player receives a favorable signal with probability 0.8 when therapy is low and with probability 0.2 when therapy is high. Signals are independent across players conditional on $T$, and each clone observes only its own signal, producing genuine imperfect information. After receiving its signal, each player chooses a phenotype from $\{P,R,Q\}$, representing proliferative (therapy-sensitive), resistant-proliferative, and quiescent (dormant) programs commonly studied in tumor ecology.
Payoffs depend on the true therapy level and the phenotypes selected: proliferation $P$ is most advantageous under low therapy, whereas resistance $R$ and quiescence $Q$ yield higher survival under high therapy. Each player selects one of these three phenotypes after observing either a favorable or unfavorable signal. The payoffs under low therapy are given by $\mathbf{A}^{\text{low}}$. We interpret these matrices as follows. If player 1 selects row $i$ and player 2 selects column $j$, then player 1 receives payoff $A^{\text{low}}_{ij}$ and player 2 receives $A^{\text{low}}_{ji}$. Similarly, the payoffs under low therapy are given by $\mathbf{A}^{\text{high}}$. Note that these differ from the final sequence-form matrix $\mathbf{A}$ and are presented to simplify the description of the game. While this game is relatively small it captures key biological features such as environmental uncertainty, private information, and phenotypic plasticity. To give an idea of the game’s size and sequential structure the full game tree is given in Figure 1.
$\mathbf{A}^{\text{low}}=\begin{array}[]{c|ccc}&P&R&Q\\ \hline\cr P&0.8&0.8&0.8\\ R&0.4&0.5&0.5\\ Q&0.4&0.5&0.5\end{array}\;\mathbf{A}^{\text{high}}=\begin{array}[]{c|ccc}&P&R&Q\\ \hline\cr P&0.1&0.1&0.1\\ R&0.1&0.6&0.2\\ Q&0.1&0.2&0.6\end{array}$
Figure 1: Extensive-form tree structure of the cancer signaling game: Nature selects therapy intensity, players receive private noisy signals and then choose phenotypes at their information sets.
$S$ AA Mean Time (s) Median Time (s) Mean #SNE (total) Mean #ESS (total) Frac. $\geq 1$ ESS
2 3 0.241 0.098 2.40 (60) 1.20 (30) 0.80
2 4 0.368 0.246 2.64 (66) 1.36 (34) 0.84
2 5 0.491 0.388 3.08 (77) 1.32 (33) 0.76
2 6 4.152 1.990 8.76 (219) 2.32 (58) 0.80
3 3 0.360 0.147 2.16 (54) 1.12 (28) 0.76
3 4 1.384 0.737 5.24 (131) 1.44 (36) 0.76
3 5 1.448 1.115 3.92 (98) 1.16 (29) 0.64
3 6 10.855 4.194 8.40 (210) 1.64 (41) 0.76
4 3 0.806 0.359 3.64 (91) 1.60 (40) 0.68
4 4 4.243 0.699 3.04 (76) 1.00 (25) 0.72
4 5 8.964 5.311 6.24 (156) 1.28 (32) 0.68
4 6 61.798 10.985 10.52 (263) 1.04 (26) 0.64
5 3 3.791 0.676 2.52 (63) 0.76 (19) 0.52
5 4 4.106 0.943 3.44 (86) 0.92 (23) 0.64
5 5 14.521 5.338 4.52 (113) 0.76 (19) 0.44
5 6 165.155 21.145 11.56 (289) 0.96 (24) 0.44
Table 2: Results for random symmetric imperfect-information extensive-form games ($K=2$ hidden states, 25 instances per setting).
Our algorithm identifies seven symmetric Nash equilibria. In every equilibrium, players assign probability 1 to the proliferative phenotype P at the “fav” information set, which reflects that proliferation is optimal when the signal indicates likely low therapy. The equilibria differ only in their behavior at the “unfav” information set. Here the QCP in Algorithm 3 uncovers a mixture of pure and mixed SNE that distribute probability across $\{P,R,Q\}$ following an unfavorable signal. Several of these arise from degeneracy: under high therapy, the resistant R and quiescent Q phenotypes yield nearly identical continuation values, producing a continuum of payoff-equivalent best responses. Table 3 summarizes all seven SNE. The full computation required only 0.287 seconds.
Unfavorable phenotype SNE ESS
$P$
$R$
$Q$
$0.5R+0.5Q$ $\times$
$0.592P+0.204R+0.204Q$ $\times$
$0.753P+0.247R$ $\times$
$0.753P+0.247Q$ $\times$
Table 3: SNE and ESS in cancer signaling game. All SNE play $P$ at the “fav” information set and differ only in their response to “unfav”.
Applying the ESS test (Algorithm 4), only three SNE are evolutionarily stable. These are exactly the realization plans that (i) play $P$ after a favorable signal and (ii) commit to a single pure phenotype $\{P,R,Q\}$ after an unfavorable signal. All mixed SNE fail the ESS test: because they equalize payoffs across the phenotypes they randomize over, a mutant can slightly shift probability mass toward the best-performing phenotype and obtain a weakly higher payoff, violating condition (2) of Definition 4. Biologically, the model therefore supports exactly three robust clonal response programs: always proliferate; proliferate on favorable signals and switch to resistant on unfavorable signals; or proliferate on favorable signals and switch to quiescent dormancy on unfavorable signals. In contrast, fine-grained mixtures among $\{P,R,Q\}$ are not evolutionarily stable and are susceptible to invasion.
To evaluate algorithm’s scalability we generalized the cancer signaling game by increasing the number of private signals $S$ and actions $A$ for each player, keeping the number of hidden states $K$ fixed at 2. For each setting of $S\in\{2,3,4,5\}$ and $A\in\{3,4,5,6\}$ we generated 25 random game instances. These games all have the same sequential structure as the original cancer signaling game, which is encoded in the sequence-form matrix $\mathbf{E}$ and vector $\mathbf{e}.$ (The cancer signaling game has $S=2$, $A=3$.) We set each entry of the $\mathbf{A}^{\text{high}}$ and $\mathbf{A}^{\text{low}}$ matrices to be uniform random in [-1,1]. We then create the sequence form payoff matrix $\mathbf{A}$ from these matrices. To reduce the number of degenerate games with a continuum of SNE we add a small amount of noise uniform in $\varepsilon_{N},\varepsilon_{N}]$ to entries of $\mathbf{A}$, using $\varepsilon_{N}=10^{-4}.$ These games form a family of random symmetric extensive-form games of imperfect information with perfect recall parametrized by $A$ and $S.$
The results of our experiments are given in Table 2. For each combination of $(S,A)$, we report the mean and median running times as well as the number of SNE found, the number of ESSs found, and the number of game instances which contained at least one ESS. For the largest games class $S=5,$ $A=6$, the median running time was 21.1 seconds, which demonstrates that the algorithm can solve moderate-sized games quickly on a laptop. We note that the runtimes were heavily skewed to the right, particularly for the larger games. This is likely because of the existence of a small number of degenerate games with a continuum of SNE. We do not report confidence intervals for mean runtimes because they are not informative for highly-skewed data, but we include the median times as well to give an idea of typical behavior.
The results in Table 2 reveal several clear trends in the structure of symmetric equilibria in random imperfect-information games. As the number of signals and actions increases, the number of SNE grows rapidly, reflecting the combinatorial expansion of possible supports and the prevalence of near-degenerate best-response structures. In contrast, ESSs remain relatively rare: even for the largest game class, only a small fraction of SNE satisfy the evolutionary stability conditions. This gap between the abundance of SNE and the scarcity of ESSs reinforces the intuition that evolutionary stability is a much stronger and more selective refinement in extensive-form settings, just as in classical normal-form models. These findings suggest that while symmetric Nash equilibria proliferate in high-dimensional information environments, genuinely stable outcomes—those robust to structured mutations—are far more constrained, underscoring the value of an exact ESS computation method.
5 Conclusion
We presented a computational framework for identifying evolutionarily stable strategies in two-player symmetric extensive-form games. The approach encodes the game exactly in sequence form, characterizes symmetric Nash equilibria through a feasibility program, and then evaluates each candidate via a nonconvex quadratic program that searches for profitable mutant deviations subject to a minimum separation constraint. Although solutions are obtained numerically and therefore depend on solver tolerances, the formulation itself captures the exact ESS and SNE conditions without requiring additional assumptions beyond perfect recall.
Our experiments demonstrate that the method is practical even for moderately sized imperfect-information games. In a biologically motivated cancer signaling game, the algorithm recovered all SNE and identified a small, interpretable subset of ESS corresponding to stable therapeutic signaling phenotypes. In the random game experiments, which span a range of signal and action complexities, we found that the number of SNE tends to grow with game size while ESS remain comparatively rare. Importantly, the computational burden was modest for most instances: median runtimes were typically well under a second for small and medium game sizes, and even the largest classes (with up to S=5S{=}5 signals and A=6A{=}6 actions) could usually be solved within a few seconds on a standard laptop. Only occasional instances exhibited substantially longer runtimes, reflecting inherent numerical difficulty rather than systematic failure of the approach.
This work suggests several promising directions for future research. On the theoretical side, it would be useful to investigate structural properties of extensive-form games that govern the number or stability of SNE, or that predict when ESS are likely to exist. On the computational side, improvements in nonlinear optimization or tailored branching strategies may further accelerate the mutant-resistance check. Biologically, the framework opens opportunities to model and analyze evolutionary stability in settings such as tumor–immune interactions, cooperative and competitive signaling among heterogeneous cancer cell subpopulations, treatment-driven evolutionary bottlenecks, and multi-phenotype ecological games where strategic behavior unfolds over latent or partially observed states. Integrating such models with empirical data may ultimately support the design of therapies that exploit evolutionary vulnerabilities in cancer ecosystems.
Our algorithm is an anytime algorithm and be halted early to output the current set of ESSs (or SNE) that have been discovered. The algorithm is sound and is guaranteed to output a subset of the full set of ESSs, though it may fail to output all ESSs if the game is degenerate and contains an infinite continuum of SNE. While we used a specific termination criterion in our experiments of halting the algorithm when no additional SNE exist with L2 distance from discovered SNE exceeding a certain threshold, many other termination criteria may be applied. For example, the algorithm could be halted after the first ESS is discovered, after a certain number of SNE have been found, or after a time limit is reached. Future work can explore improved techniques for addressing degeneracy.
While our algorithm was presented in the context of two-player (symmetric) games, the algorithm can be extended straightforwardly to the case of $n>2$ players using techniques similar to a recent approach for Nash equilibrium computation Ganzfried (2025b). The definition of multiplayer ESS can be extended straightforwardly to the extensive-form setting, and the concepts of symmetric games and symmetric Nash equilibria can be defined analogously. The main difference is that the equivalent version of the $\mathbf{A}\mathbf{x}$ terms in Algorithms 2 and 3 for finding the next SNE will now contain products of n1n-1 variables. We can still represent these constraints as quadratic but will need to introduce a number of auxiliary variables that is exponential in n1.n-1. For small games with 3 or 4 players this approach may be feasible, but new algorithmic advances would most likely be needed to solve larger games. As for the normal-form case, the QCQP for testing whether an SNE is an ESS would not be changed if the number of players is increased Ganzfried (2025a).
Our algorithm is fundamentally different from approaches for computing ESS in normal-form games. Those approaches are based on enumerating supports, and for each support computing a potential SNE and ESS. While the number of supports is exponential in the number of pure strategies, for relatively small games it can be done efficiently in practice. By contrast, it does not make sense to think of the concept of supports in extensive-form games, as enumeration of all of them is infeasible. We also note that several efficient preprocessing procedures developed for the normal-form setting are no longer applicable to the extensive form. A recent approach applies a “strict NE shortcut” procedure that immediately classifies an SNE as an ESS if it satisfies a simple property ensuring that it is a strict equilibrium Ganzfried (2025a). That approach involves testing whether there is a unique best response to the SNE, which is simple to test in normal-form games but not in extensive-form games. That work also applies a procedure that discards SNE that fail a “pure mutant screen,” which is also easy to do in normal-form games but not in extensive-form games. By contrast our approach creates a QCPC that searches for potential pure and mixed mutants. Perhaps improved preprocessing procedures can be developed that are applicable to extensive-form games and can reduce the number of QCPC solves needed by Algorithm 4.